Consider the following table of standard electrode potentials for a series of hypothetical reactions in aqueous solution:

Reduction Half-Reaction E˚ (v)
X to the power of plus open parentheses a q close parentheses space plus space e to the power of minus rightwards arrow space straight X open parentheses straight s close parentheses 1.54
Y to the power of 2 plus end exponent open parentheses a q close parentheses space plus space 2 space e to the power of minus rightwards arrow space straight Y open parentheses straight s close parentheses 1.16
Z to the power of plus open parentheses a q close parentheses space plus space e to the power of minus rightwards arrow space straight Z open parentheses straight s close parentheses 0.84

Calculate the standard free-energy change, ∆G˚, in units of kJ/mol at 298 K for the following reaction

Y to the power of 2 plus end exponent open parentheses a q close parentheses space plus space 2 space straight Z open parentheses s close parentheses space rightwards arrow space straight Y open parentheses straight s close parentheses space plus space 2 space straight Z to the power of plus open parentheses aq close parentheses
Answer #1

Oxidation reaction : 2 Z (s) \rightarrow 2 Z + (aq) + 2 e -

Reduction reaction : Y 2+ (aq) + 2 e -\rightarrow Y (s)  

Overall reaction : 2 Z (s) + Y 2+ (aq) \rightarrow 2 Z + (aq) + Y (s)  

In overall reaction , two electrons are transferred.

The standard emf of the cell is calculated as E 0 cell = E 0 cathode - E 0 anode

\therefore E 0 cell = E 0 Y 2+ I Y - E 0 Z + I Z

\therefore E 0 cell = 1.16 V - 0.84 V = 0.32 V

We have relation between standard Gibbs energy change and emf of the cell as \Delta G 0 = - n F E 0 cell

\therefore\Delta G 0 = - 2 \times 96487 C/ mol \times 0.32 V

\therefore\Delta G 0 = - 61751.7 J /mol

\therefore\Delta G 0 = - 61.75 kJ / mol

ANSWER : \Delta G 0 = - 61.75 kJ / mol